OpenStudy (anonymous):

what does this symbol represent, I just finished studying limits of finite sums

OpenStudy (anonymous):

\[\sum_{}^{}n/k=1 f(c_k)(1/n)\]

OpenStudy (anonymous):

whats does the n stand for

OpenStudy (anonymous):

what about k

OpenStudy (anonymous):

at least it tend to

OpenStudy (anonymous):

i though that would be like te number of sub intervals

myininaya (myininaya):

\[\sum_{k=1}^{n}\frac{n}{k}=n \sum_{k=1}^{n}\frac{1}{k}=n(\frac{1}{1}+\frac{1}{2}+\frac{1}{3}+....+\frac{1}{n})\]

myininaya (myininaya):

i don't know if you mean for the k to be changing or the n i just assumed the k

myininaya (myininaya):

\[\sum_{n=1}^{k}\frac{n}{k}\]?

myininaya (myininaya):

which do you mean

OpenStudy (anonymous):

This is what i mean